# Alternative ranking measures to predict international football results

### v1

@@ Abstract
Over the last few years, there has been a growing interest in the prediction and modelling of competitive sports outcomes, with particular emphasis placed on this area by the Bayesian statistics and machine learning communities. In this paper, we have carried out a comparative evaluation of statistical and machine learning models to assess their predictive performance for the 2022 World Cup and for the 2024 Africa Cup of Nations by evaluating alternative summaries of past performances related to the involved teams. More specifically, we consider the Bayesian Bradley-Terry-Davidson model, which is a widely used statistical framework for ranking items based on paired comparisons that have been applied successfully in various domains, including football. The analysis was performed including in some canonical goal-based models both the Bradley-Terry-Davidson derived ranking and the widely recognized Coca-Cola FIFA ranking commonly adopted by football fans and amateurs.
@@ 1 Introduction
This paper aims to investigate the potential improvement in predictive performance achieved by some statistical goal-based methods and machine learning result-based algorithms when an appropriate ranking is used as an additional predictor. To asses it, we analyze the outcomes of the the most recent 2022 FIFA World Cup in Qatar and the 2024 Africa Cup in Ivory Coast. We considered a Bayesian Bradley-Terry-Davidson derived ranking system, using the posterior median of the log-strength parameters as novel predictor. Subsequently, we compare the predictive performances of this approach with those obtained using the well-established FIFA ranking to determine which of them provides more reliable forecasts. The rest of the paper is organized as follows. Section 2 delves into the theoretical framework, introducing the standard Bradley-Terry model and its Davidson extension for handling draws, followed by a discussion of the Bayesian approach. Furthermore, Section 3 describes the statistical goal-based methods and the machine learning result-based algorithms used in this study. In Section 4, we evaluate the application of these methodologies on the data from both the last World Cup and Africa Cup of Nations. Finally, Section 5 provides concluding remarks, outlining the limitations, advantages, and potential future research directions.
@@ 3.1 Goal-based models
Goal-based models assume that the number of goals scored in a game by each team follows a discrete distribution, typically two independent Poisson or a bivariate-Poisson accounting for positive correlation. Thus, for each game, we need to consider the pair of counts $(X_{in},Y_{jn})$, for $i\neq j=1,\dots,N_{T}$ and $n=1,\dots,N$. The first count $X_{in}$ denotes the non-negative number of goals scored by the home team $i$ and the second count $Y_{jn}$ denotes the number of goals scored by the visiting team j, both in the n-th game. A simple double Poisson model is
@@ 4.1 FIFA 2022 World Cup in Qatar
Figure 1: World Cup 2022: the top panel displays scatterplots comparing the ranks of teams in both the group stage (orange dots) and knockout stage (green dots) under the two ranking systems. The dashed grey line represents the bisector. The bottom panel presents boxplots of rank differences for the FIFA ranking (blue) and the Bayesian BTD derived ranking (red) across the two World Cup stages.
@@
Figure 2: World Cup 2022: Brier score for the FIFA ranking (blue) and the Bayesian BTD derived ranking (red) across the two World Cup stages.
4.2 2024 Africa Cup of Nations
In this section, we fit the considered statistical models and machine learning algorithms to the data from the most recent Africa Cup of Nations (AFCON) tournament held in Ivory Coast. Notably, the AFCON competition differs from the FIFA World Cup in that the participating teams tend to be more similar in terms of overall strength even during the group stage, making it a compelling case for analyzing the effectiveness of the Bayesian BTD ranking system. In order to conduct this analysis, we use as training set the data from matches played throughout the period 2018 to the end of 2023. Furthermore, the Bayesian BTD model was executed within this same period to generate team rankings. In addition, the FIFA ranking employed correspond to those published on December 21st, available at https://inside.fifa.com/fifa-world-ranking/men?dateId=id14233.
@@
Figure 3: Africa Cup 2023: the top panel displays scatterplots comparing the ranks of teams in both the group stage (orange dots) and knockout stage (green dots) under the two ranking systems. The dashed grey line represents the bisector. The bottom panel presents boxplots of rank differences for the FIFA ranking (blue) and the Bayesian BTD derived ranking (red) across the two Africa Cup stages.
@@
Figure 4: Africa Cup 2023: Brier score for the FIFA ranking (blue) and the Bayesian BTD derived ranking (red) across the two Africa Cup stages.
@@ 5 Discussion
This paper investigates the potential improvement in predictive performance of statistical goal-based methods and machine learning result-based algorithms when a ranking system is incorporated as an additional predictor variable. We analyze data from the recent 2022 FIFA World Cup in Qatar and the 2023 Africa Cup of Nations in Ivory Coast. Specifically, we explore the effectiveness of a Bayesian Bradley-Terry-Davidson derived ranking system in enhancing prediction accuracy compared to the well-established FIFA ranking system. We compare the performance of these two ranking systems across different tournament stages to identify their relative strengths and weaknesses in predicting match outcomes.

### v2

@@ Abstract
Over the last few years, there has been a growing interest in the prediction and modelling of competitive sports outcomes, with particular emphasis placed on this area by the Bayesian statistics and machine learning communities. In this paper, we have carried out a comparative evaluation of statistical and machine learning models to assess their predictive performance for the 2022 FIFA World Cup and for the 2023 CAF Africa Cup of Nations by evaluating alternative summaries of past performances related to the involved teams. More specifically, we consider the Bayesian Bradley-Terry-Davidson model, which is a widely used statistical framework for ranking items based on paired comparisons that have been applied successfully in various domains, including football. The analysis was performed including in some canonical goal-based models both the Bradley-Terry-Davidson derived ranking and the widely recognized Coca-Cola FIFA ranking commonly adopted by football fans and amateurs.
@@ 1 Introduction
This paper aims to investigate the potential improvement in predictive performance achieved by some statistical goal-based methods and machine learning result-based algorithms when an appropriate ranking is used as an additional predictor. To asses it, we analyze the outcomes of the the most recent 2022 FIFA World Cup in Qatar and the 2023 Africa Cup in Ivory Coast. We considered a Bayesian Bradley-Terry-Davidson derived ranking system, using the posterior median of the log-strength parameters as novel predictor. Subsequently, we compare the predictive performances of this approach with those obtained using the well-established FIFA ranking to determine which of them provides more reliable forecasts. The rest of the paper is organized as follows. Section 2 delves into the theoretical framework, introducing the standard Bradley-Terry model and its Davidson extension for handling draws, followed by a discussion of the Bayesian approach. Furthermore, Section 3 describes the statistical goal-based methods and the machine learning result-based algorithms used in this study. In Section 4, we evaluate the application of these methodologies on the data from both the last World Cup and Africa Cup of Nations. Finally, Section 5 provides concluding remarks, outlining the limitations, advantages, and potential future research directions.
@@ 3.1 Goal-based models
Goal-based models assume that the number of goals scored in a game by each team follows a discrete distribution, typically two independent Poisson or a bivariate-Poisson accounting for positive correlation. Thus, for each game, we need to consider the pair of counts $(X_{in},Y_{jn})$, for $i\neq j=1,\dots,N_{T}$ and $n=1,\dots,N$. The first count $X_{in}$ denotes the non-negative number of goals scored by the home team $T_{i}$ and the second count $Y_{jn}$ denotes the number of goals scored by the visiting team $T_{j}$, both in the n-th game. A simple double Poisson model is
@@ 4 Applications
Figure 1: 2022 FIFA World Cup. The top panel displays scatterplots comparing the ranks of teams in both the group stage (orange dots) and knockout stage (green dots) under the two ranking systems. The dashed grey line represents the bisector. The bottom panel presents boxplots of rank differences for the FIFA ranking (blue) and the Bayesian BTD derived ranking (red) across the two World Cup stages.
@@
Figure 2: 2022 FIFA World Cup. Brier score for the FIFA ranking (blue) and the Bayesian BTD derived ranking (red) across the two World Cup stages.
4.2 2023 CAF Africa Cup of Nations
In this section, we fit the considered statistical models and machine learning algorithms to the data from the most recent CAF Africa Cup of Nations (AFCON) tournament held in Ivory Coast. Notably, the AFCON competition differs from the FIFA World Cup in that the participating teams tend to be more similar in terms of overall strength even during the group stage, making it a compelling case for analyzing the effectiveness of the Bayesian BTD ranking system. In order to conduct this analysis, we use as training set the data from matches played throughout the period 2018 to the end of 2023. Furthermore, the Bayesian BTD model was executed within this same period to generate team rankings. In addition, the FIFA ranking employed correspond to those published on December 21st, available at https://inside.fifa.com/fifa-world-ranking/men?dateId=id14233.
@@
Figure 3: 2023 CAF Africa Cup of Nations. The top panel displays scatterplots comparing the ranks of teams in both the group stage (orange dots) and knockout stage (green dots) under the two ranking systems. The dashed grey line represents the bisector. The bottom panel presents boxplots of rank differences for the FIFA ranking (blue) and the Bayesian BTD derived ranking (red) across the two Africa Cup of Nations stages.
@@
Figure 4: 2023 CAF Africa Cup of Nations. Brier score for the FIFA ranking (blue) and the Bayesian BTD derived ranking (red) across the two Africa Cup of Nations stages.
@@ 5 Discussion
This paper investigates the potential improvement in predictive performance of statistical goal-based methods and machine learning result-based algorithms when a ranking system is incorporated as an additional predictor variable. We analyze data from the recent 2022 FIFA World Cup in Qatar and the 2023 CAF Africa Cup of Nations in Ivory Coast. Specifically, we explore the effectiveness of a Bayesian Bradley-Terry-Davidson derived ranking system in enhancing prediction accuracy compared to the well-established FIFA ranking system. We compare the performance of these two ranking systems across different tournament stages to identify their relative strengths and weaknesses in predicting match outcomes.